OpenStudy (anonymous):

Find the center, vertices, foci, and the equations of the asymptotes of the hyperbola: x^2+9y^2+36y-72=0

OpenStudy (mertsj):

Put the 72 on the right side.

OpenStudy (mertsj):

Then write \[9y^2+36y=9(y^2+4y+\]

OpenStudy (mertsj):

And complete the square.

OpenStudy (anonymous):

so ignore the x^2?

OpenStudy (mertsj):

Only for a moment.

OpenStudy (mertsj):

Just until we get that y slop taken care of.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

so is it x^2+9(y+2)^2=108?

OpenStudy (mertsj):

Wow!! I am impressed. Yes. That's it. Now divide every term by 108

OpenStudy (mertsj):

Oh. Something is wrong. This is not a hyperbola. Is there a typo somewhere?

OpenStudy (anonymous):

my teacher said it is a hyperbola so.

OpenStudy (mertsj):

Are you sure there is not a sign error? In hyperbolas, either x2 or y^2 is negative.

OpenStudy (mertsj):

If they are both positive, it is either an ellipse or a circle.

OpenStudy (anonymous):

whoops there is a sign error its supposed to be a -9y^2. sorry, my bad.

OpenStudy (mertsj):

So we need to rethink that y part because now it is: -9(y^2-4y

OpenStudy (anonymous):

yeah you're right. so it would be like x^2-9(y-2)^2=36?

OpenStudy (mertsj):

Yes. That's it. Now divide each term by 36

OpenStudy (anonymous):

okay so it'd be \[\frac{ x^2 }{ (6)^2 }- \frac{ (y-2)^2 }{ (2)^2 }=1 \]

OpenStudy (mertsj):

Yes. So what is the center?

OpenStudy (anonymous):

(0,2)

OpenStudy (mertsj):

And what is the value of a?

OpenStudy (anonymous):

is it 0?

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]

OpenStudy (mertsj):

Your equation matches that one. h=0, k = 2, what is a?

OpenStudy (anonymous):

it would equal 6.

OpenStudy (mertsj):

Since x^2 is positive, this hyperbola opens to the right and to the left. And the vertices are 6 units to the right and the left of the center.

OpenStudy (mertsj):

So what are the vertices?

OpenStudy (anonymous):

okay. so it'd be (-6,2) and (6,2)

OpenStudy (mertsj):

Yes. Now we should probably put this information on a graph.

OpenStudy (mertsj):



OpenStudy (mertsj):

Did your teacher teach you to draw the rectangle based on the a and b values?

OpenStudy (anonymous):

nope not really. he just showed us the math to get the information. he sorta skipped over it.

OpenStudy (mertsj):

a=6 and b = 2. So if we go a units right and left of the center, and b units up and down from the center we can draw a rectangle based on those points.

OpenStudy (mertsj):